OpenStudy (lukecrayonz):

Find the nth term of the geometric sequence: a_2 or a[2]=3, a_5 or a[5]=(3/64), n=1

OpenStudy (lukecrayonz):

@satellite73

OpenStudy (lukecrayonz):

@bahrom7893 anything? haha

OpenStudy (lukecrayonz):

I literally know the equation if you have a[1], but nothing when you're solving for a[1]

OpenStudy (lukecrayonz):

http://screensnapr.com/v/3ZOlWC.png maybe this will help

OpenStudy (lukecrayonz):

Because the equation while using a[1], is a[n]=a[1]r^(n-1)

OpenStudy (bahrom7893):

a2=a1*r=3; a1=3/r

OpenStudy (bahrom7893):

just do the same thing as the solved example.. or is there a trick somewhere?

OpenStudy (lukecrayonz):

And solving for r now.. other example doesn't make that much sense to me, and I feel like there is a trick because n-1, 1-1=0.

OpenStudy (lukecrayonz):

and anything to the power of 0 is 1, so a[1]*1=a[1]..?

OpenStudy (bahrom7893):

r is either powers of 4 or 2..

OpenStudy (lukecrayonz):

Hmm..?

OpenStudy (bahrom7893):

hang on..

OpenStudy (bahrom7893):

r=1/4, a1=3/1/4=12

OpenStudy (bahrom7893):

a2=3=a1r a1=3/r

OpenStudy (bahrom7893):

a5=a1r^4=a1rr^3=3r^3=3/64

OpenStudy (bahrom7893):

r^3=1/64 r=1/4

OpenStudy (bahrom7893):

a1=3/r=3/1/4=12

OpenStudy (lukecrayonz):

Got it! Thank you so much haha, stumped me

OpenStudy (bahrom7893):

i don't even know why they give u n, it's totally irrelevant imo

OpenStudy (lukecrayonz):

Got another question, that literally has nothing to do with anything..

OpenStudy (lukecrayonz):

Posted it